OpenStudy (anonymous):

Calc I Problem: Need some help with these problems, they are test problems that i think i got correct and were marked wrong. Find the equation of the tangent line to the graph of f(x)=x^2 at the point (2,4). A.) Find the value of f'(2). You may use the rule in order to find the derivative. B.) So, the slope of the tangent line is _____. C.) The equation of the tangent line is f(x) = ____.

OpenStudy (phi):

what did you answer?

OpenStudy (anonymous):

My answers are: A) f'(2) = 4 B) 2 C) 2x

OpenStudy (jonnyvonny):

Monitoring.

OpenStudy (anonymous):

oooh..damn so thats why i missed it. but the slope = 2 is correct?

OpenStudy (phi):

wait, y=2x does work. y= 2*2 = 4 so you are correct.

OpenStudy (anonymous):

so the whole answer is correct?

OpenStudy (phi):

you found the correct f'(2) = 4 how did you get the slope = 2 ( the derivative is the slope)

OpenStudy (jonnyvonny):

Well, its asking for the slope when x=2, so although you got the derivative of the equation correct, you did not get the slope when x=2

OpenStudy (anonymous):

ooh i just figured that y=2x is the equation so the slope was 2 'y=mx+b'

OpenStudy (anonymous):

m=2

OpenStudy (phi):

there are lots of lines that could go through (2,4) you want the line that has the *same* slope as the tangent line i.e. the derivative's value (at x=2)

OpenStudy (anonymous):

so the slope would have been 4? :\

OpenStudy (phi):

yes. The derivative is the "slope" at a *very short section" of the curve

OpenStudy (anonymous):

damn alright so thats why i got that wrong.

OpenStudy (phi):

once you use m=4 you have to find b to answer the last part.

OpenStudy (anonymous):

well she gave me partial credit so i guess thats pretty fair. i had it just gave the wrong part for the answer.

OpenStudy (anonymous):

alright well that makes sense now, thank you.

OpenStudy (phi):

yw